The molality of 0.05 mol hcl in 5 litre solution at 4°c


1
Expert's answer
2021-06-07T03:15:30-0400

Let the volume of the solution be 1 litre

Molarity = moles of NaOH/volume of the solution in litres

0.05 moles of NaOH

Mass of solvent = Volume × Density

= 1×1

= 1Kg

Molar Mass of NaOH = 39.997 × 0.05

= 1.99985

= 1.99985/1

= 1.99985